# The topology of a chaotic attractor in the Kuramoto-Sivashinsky equation

### v1

@@ V Conclusion
In this paper, we have found a candidate template for the topology of a chaotic attractor in the Kuramoto-Sivashinski equation. To our knowledge, this is the first attempt to do this in a PDE. While it is possible that our template would be shown to be invalid when other periodic orbits are included, it was found using only five UPOs and validated with a further ten, so we are confident it gives an accurate representation of the topology of the dynamics. Furthermore, equivalent topologies were found using two different dimensionality-reduction techniques: proper orthogonal decomposition and an autoencoder. This increases our confidence of the accuracy of the topology, and furthermore provides strong evidence that these dimensionality reduction methods are sufficiently accurate to capture the behaviour of the system in this case.
The main result of this paper is figure 13 (or equivalently figure 14). We now have a qualitative interpretation of the dynamics: as trajectories move around the chaotic attractor, they are divided into two branches, which intertwine and eventually merge. The bounding torus for this attractor has one hole, unlike the famous Lorenz attractor, but like the classical ODE systems of the Rössler and Sprott D attractors (Letellier, Stankevich, and Rössler, 2022). As with the templates for those latter ODE systems, the KSE attractor exhibits two branches, one of which has an additional twist, which demonstrates stretching-and-folding chaos (as opposed to ‘tearing-and-squeezing’, the other standard mechanism). Compared with those ODE systems however, the KSE attractor has many more braids.
Topological descriptions of chaotic dynamics give a way of quantifying the stretching and folding processes which give rise to chaos. In a three-dimensional system like the Lorenz model, this folding is directly visible from looking at the state space of the attractor. In higher-dimensions, projections and dimensionality reduction methods do not immediately provide this information, but a diagram of the template, as given in figures 13 and 14, provide an immediate and intuitive interpretation of the dynamics. A topological description of a chaotic dissipative PDE is particularly useful, as it can be used to compare between different dimensionality reduction techniques, and different discretisations of the system, to confirm that the essential dynamics are preserved, even if the exact quantitive details of the system may have been distorted.
@@
However, hope for applying topological methods to other PDEs is given by more modern techniques: assigning an $n$-dimensional cell-complex to a cloud of points on the attractor allows one to calculate invariants such homology groups, Betti number and Euler characteristic which provide information about the topological structure (Muldoon et al., 1993; Lefranc, 2013; Sciamarella and Mindlin, 1999, 2001). BraMAH (branched manifold analysis through homologies) can be applied to identify torsions, branch crossings and weak boundaries (Charó, Artana, and Sciamarella, 2021; Charó et al., 2020). Recent work has improved this through the notion of templex (Charó, Letellier, and Sciamarella, 2022) by endowing it with an oriented graph encoding the direction of the flow. A templex associated with a four-dimensional system is constructed in Charó, Letellier, and Sciamarella (2022). Though unlikely to be successful in the full state space of the discretised PDE, given the computational complexity of such a high-dimensional system, this is certainly possible in 4- or 5-dimensions, and so our approach of using an autoencoder to reduce the dimension can be combined with templexes to study the topology of more complicated PDE systems.
@@ Acknowledgements.
The authors wish to thank Omid Ashtari for fruitful discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant no. 865677).

### v2

@@ V Conclusion
In this paper, we have found a candidate template for the topology of a chaotic attractor in the Kuramoto-Sivashinski equation. To our knowledge, this is the first attempt to do this in a PDE. While it is possible that our template would be shown to be invalid when other periodic orbits are included, it was found using only five UPOs and validated with a further ten, so we are confident it gives an accurate representation of the topology of the dynamics. Furthermore, equivalent topologies were found using two different dimensionality-reduction techniques: proper orthogonal decomposition and an autoencoder. Though neither method is guaranteed to preserve the topology in general (Uribarri and Mindlin, 2020), the agreement between the two increases our confidence of the accuracy of the topology, and furthermore provides strong evidence that these dimensionality reduction methods are sufficiently accurate to capture the behaviour of the system in this case.
The main result of this paper is figure 13 (or equivalently figure 14). We now have a qualitative interpretation of the dynamics: as trajectories move around the chaotic attractor, they are divided into two branches, which intertwine and eventually merge. The bounding torus for this attractor has one hole, unlike the famous Lorenz attractor, but like the classical ODE systems of the Rössler and Sprott D attractors (Letellier, Stankevich, and Rössler, 2022). As with the templates for those latter ODE systems, the KSE attractor exhibits two branches, one of which has an additional twist, which demonstrates stretching-and-folding chaos (as opposed to ‘tearing-and-squeezing’, the other standard mechanism).
Topological descriptions of chaotic dynamics give a way of quantifying the stretching and folding processes which give rise to chaos. In a three-dimensional system like the Rössler system, this folding is directly visible from looking at the state space of the attractor. In higher-dimensions, projections and dimensionality reduction methods do not immediately provide this information, but a diagram of the template, as given in figures 13 and 14, provide an immediate and intuitive interpretation of the dynamics. A topological description of a chaotic dissipative PDE is particularly useful, as it can be used to compare between different dimensionality reduction techniques, and different discretizations of the system, to confirm that the essential dynamics are preserved, even if the exact quantitive details of the system may have been distorted.
@@
However, hope for applying topological methods to other PDEs is given by more modern techniques: assigning an $n$-dimensional cell-complex to a cloud of points on the attractor allows one to calculate invariants such homology groups, Betti number and Euler characteristic which provide information about the topological structure (Muldoon et al., 1993; Lefranc, 2013; Sciamarella and Mindlin, 1999, 2001). BraMAH (branched manifold analysis through homologies) can be applied to identify torsions, branch crossings and weak boundaries (Charó, Artana, and Sciamarella, 2021; Charó et al., 2020). Recent work has improved this through the notion of templex (Charó, Letellier, and Sciamarella, 2022) by endowing it with an oriented graph encoding the direction of the flow. A templex associated with a four-dimensional system is constructed in Charó, Letellier, and Sciamarella (2022). Though unlikely to be successful in the full state space of the discretized PDE, given the computational complexity of such a high-dimensional system, this is certainly possible in 4- or 5-dimensions, and so our approach of using an autoencoder to reduce the dimension can be combined with templexes to study the topology of more complicated PDE systems.
@@ Acknowledgements.
The authors are particularly indebted to an anonymous referee for finding an error in an initial version of the graphical representation of the template at the heart of this paper. The authors wish to thank Omid Ashtari for fruitful discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant no. 865677).